A campus student club distributed material about membership to new students attending an orientation meeting. Of those receiving this material 40% were men and 60% were women. Subsequently, it was found that 7% of the men and 9% of the women who received this material joined the club.
a. Find the probability that a randomly chosen new
student who receives the membership material will join the club.
b. Find the probability that a randomly chosen new student who joins the club after receiving the membership material is a woman.
1
Expert's answer
2020-01-29T16:53:58-0500

a.

"P(men)=0.4, \ P(women)=0.6,"

"P(join|men)=0.07, \ P(join|women)=0.09"


"P(join)=P(men)P(join|men)+P(women)P(join|women)"

"P(join)=0.4(0.07)+0.6(0.09)=0.082"

b.


"P(women|join)={P(women)P(join|women) \over P(join)}"

"P(women|join)={0.6(0.09) \over 0.082}={27 \over 41}\approx0.659"